OpenStudy (anonymous):

Find the area of the region that is surrounded by the curves \(y = -3 + 11 x - x^2\) and \(y = -3 + 4 x\)

OpenStudy (anonymous):

ANyone?

OpenStudy (anonymous):

yeah, do you have preference for any integral methods ? like Single Variable integrals or do you want to use double integrals. both ways you get the same answer

OpenStudy (anonymous):

um... whatever is easier

OpenStudy (anonymous):

well both are easy ways, but it depends on you, are you familiiar with double integrals

OpenStudy (anonymous):

Um not really I dont really know where to star just do wht we would learn in a second calsulus class

OpenStudy (anonymous):

you can use Area = \[\int\limits_{a}^{b}\int\limits_{y=f(x)}^{y=g(x)}\] dA

OpenStudy (anonymous):

where dA= dy dx , integrating with respect to y first and then with respect to x

OpenStudy (anonymous):

You lost me

OpenStudy (anonymous):

to find values of a and b , you will need to have -3+4x=-3+11x-x^2

OpenStudy (anonymous):

solve for values of x and that should give you the bounds of integral, i.e. lower limit and and upperlimit

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

Can you draw a picture?

OpenStudy (anonymous):

graphing is not easy, I mean it takes lot of space here

OpenStudy (anonymous):

ohh okay I thought I cold solve it using integrals and no pics?

OpenStudy (anonymous):

Yikes that's alot to read